OpenStudy (anonymous):

Which trig expression is equal to sin (72° - a)?

OpenStudy (accessdenied):

It seems that you were offered choices here, but there are none listed. Can you complete the problem statement so that we may assist you properly? My only hypothesis is that we are looking at the angle sum formula of sine: sin (a + b) = sin a cos b + sin b cos a For a=72, b = -a. sin(72 + (-a)) = sin (72 - a) = what? (substitute into the formula).